(python only) Write a program that asks the user for their filing status (single or married) and their taxable income. Then, using the below table, compute the tax owed and display the filing status, taxable income, and tax owed.
If the filing status is single and the taxable income is overbut not over   the tax is of the amount over$0$9,52510%0$9,526$38,700$952.50 + 12%$9,525$38,701$82,500$4,453.50 +22%
$38,700
$82,501unlimited$14,089.50 +24%
$82,500
If the filing status is married filing jointly and the taxable income is overbut not overthe tax is of the amount over$0$19,05010%$0$19,051$77,400$1,905 + 12%$19,050$77,401$165,000$8,907 + 22%$77,400$165,001unlimited$28,179 + 24%$165,000

Here are some check figures:


Single filer with $50,000 taxable income owes $6,939.50 tax


Married filer with $25,000 taxable income owes $2,619.00 tax


Married filer with $200,000 taxable income owes $36,579.00 tax


Be sure to follow the Program Submission Guidelines document from last week. Save the program and upload the .py file here.
Answer #1

#taking input from the user
mstat = input("Enter single or married ")
inc = float(input("Enter your income "))
#calculating tax if status is single
if mstat=='single':
if inc > 0 and inc <=9525:
tax=0.1*inc
elif inc > 9526 and inc <=38700:
tax=952.50 + 0.12*(inc-9525)
elif inc > 38701 and inc <=82500:
tax=4453.50 + 0.22*(inc-38700)
else:
tax=14089.50 + 0.24*(inc-82500)
#calculating tax if status is married
if mstat=='married':
if inc > 0 and inc <=19050:
tax=0.1*inc
elif inc > 19051 and inc <=77400:
tax=1905 + 0.12*(inc-19050)
elif inc > 77401 and inc <=16500:
tax=8907 + 0.22*(inc-77400)
else:
tax=28179 + 0.24*(inc-165000)
#displaying the result
print("Marital Status : ",mstat)
print("Income : $",inc)
print("Tax To Pay : $",tax)

OUTPUT